0 0
Add a comment Improve this question Transcribed image text
Answer #1

Answer-

Given

The likelihood of economic boom = 75 % =0.75

Therefore the likelihood of economic recession = 1 - 75 % = 1 - 0.75 = 0.25

return on stock A during economic boom = 14 %
return on stock A during economic recession = - 4 %

The expected recession on Stock A = return on stock A during economic boom x likelihood of economic boom + return on stock A during economic recession x the likelihood of economic recession

The expected recession on Stock A = 14 % x 0.75 + ( - 4 % ) x 0.25

The expected recession on Stock A = 10.5 % - 1 %

The expected recession on Stock A = 9.5 %